OpenStudy (silverfang492):

how do I simplify \(1+4+9+\cdot\cdot\cdot+x^2\)

OpenStudy (zzr0ck3r):

\(\sum_{i=1}^ni^2=1+4+9+16+..+n^2\)

OpenStudy (kainui):

Whether this is what you want or not, oh well I enjoy this too much to stop from writing it out. Take note of this telescoping series: \[\sum_{k=1}^x k^3 - (k-1)^3 = [1^3-0^3]+[2^3-1^3]+\cdots + [x^3 - (x-1)^3]= x^3\] Now you can expand the inner sum: \[\sum_{k=1}^x k^3-(k-1)^3 = x^3\]\[\sum_{k=1}^x k^3 - (k^3-3k^2+3k-1) =x^3\]\[\sum_{k=1}^x 3k^2-3k+1 = x^3\] You can pick apart this sum of 3 terms into 3 sums: \[3\sum_{k=1}^x k^2-3\sum_{k=1}^x k+\sum_{k=1}^x1 = x^3\] The first sum is what we want, the second sum you might recognize from Gauss's trick (one of his many tricks really lol) and the last sum is just counting. \[\sum_{k=1}^xk = \frac{x(x+1)}{2}\]\[\sum_{k=1}^x 1 = x\] Throw those in to get: \[3\sum_{k=1}^x k^2-3\sum_{k=1}^x k+\sum_{k=1}^x1 = x^3\] \[3\sum_{k=1}^x k^2-3\frac{x(x+1)}{2}+x = x^3\] Little algebra away now: \[\sum_{k=1}^x k^2 = \frac{x^3}{3} + \frac{x(x+1)}{2} - \frac{x}{3}\] Little more tidied up: \[\sum_{k=1}^x k^2 = \frac{x(x+1)(2x+1)}{6}\] Hopefuly I didn't make any mistakes, if I did pls forgive the idea is still in tact I hope. The point is you can easily(heh) make a telescoping series, and from that you can get an expression for any sum of powers.